OpenStudy (anonymous):

5w+6=9-w help full instructions soooo confused

jimthompson5910 (jim_thompson5910):

5w+6=9-w 5w=9-w-6 5w+w=9-6 6w=9-6 6w=3 w=3/6 w=1/2

jimthompson5910 (jim_thompson5910):

The basic idea is to get w isolated on one side and everything else to the other side

OpenStudy (anonymous):

thanks mate :)